In order to construct a fence around a rectangular field whose length exceeds twice its width by 5 meters, 160 meters of fencing are required. What is the area of the field?
Answer by Boreal(15235)

L=2W+5
L+W=80, since half perimeter is L+W
2W+5+W=80
3W=75
W=25 m
L=55 m
The perimeter is 2(55)+2(25)=160 m
The area is 55*25=1375 m^2